# Statistical learning theory and Occam’s razor: The argument from empirical risk minimization

### v1

@@ Abstract.
This paper considers the epistemic justification for a simplicity preference in inductive inference that may be obtained from the machine learning framework of statistical learning theory. Uniting elements from both earlier arguments suggesting and rejecting such a justification, the paper spells out a qualified means-ends and model-relative justificatory argument, built on statistical learning theory’s central mathematical learning guarantee for the method of empirical risk minimization.
@@ 1. Introduction
Statistical learning theory is the standard framework for the mathematical analysis of machine learning methods (Vapnik, 2000; Shalev-Shwartz and Ben-David, 2014). The framework offers theoretical learning guarantees for certain algorithmic learning methods, thus providing a foundation for holding that such methods are good methods.
An old trope in machine learning, usually invoked under the label Occam’s razor, is that a shared trait of many good methods is a bias towards simplicity (Mitchell, 1997; Duda et al., 2001; Shalev-Shwartz and Ben-David, 2014; Goodfellow et al., 2016; Alpaydin, 2020). Occam’s razor, understood as the principle that a simplicity preference is integral to good scientific or inductive reasoning, is also a long-standing topic of debate in the philosophy of science (Sober, 2015; Baker, 2022). The central question here is whether we actually have some epistemic justification for Occam’s razor. That is, we seek a rational reason for holding that a simplicity preference helps us attain desirable epistemic ends, like minimizing error.
@@
In this paper, I develop a qualified epistemic justification for Occam’s razor in statistical learning theory, which accommodates the intuitions and arguments from both sides. An important component of my account is the model-relativity of the justification obtainable from theoretical learning guarantees, as highlighted by Sterkenburg and Grünwald (2021). Another important characteristic is the means-ends nature of the justificatory reasoning. The result—with a label that is testimony to its highly qualified nature—is a means-ends model-relative justification for Occam’s razor in statistical learning theory.
@@
The argument I present here is based on the first of the two “inductive principles” (Vapnik, 2000) central to statistical learning theory, namely the method of empirical risk minimization. The argument does also inherit from this method a certain restriction in scope. However, the current argument forms, I think, the core for a more general argument, based on the more general method of structural risk minimization. But that is work for another paper.
Apart from the qualifications to the argument, there is a worry about the paper’s general project that I should acknowledge. This worry is that the project engages with a debate of a bygone era. It has been well over a decade since Harman and Kulkarni (2007) initiated a small wave of philosophical interest in statistical learning theory, and Steel (2011, p. 860) concluded that the theory is “worthy of further sustained interest from philosophers of science.” This sustained interest has not exactly materialized, while the landscape of machine learning has altered significantly. Especially the advent of deep neural networks (DNN’s) has caused a shift in what seem the more pertinent epistemological issues: from the traditional questions around the reliability of inductive inference to questions around interpretability and explainability (Beisbart and Räz, 2022; Freiesleben and König, forthcoming). Moreover, the advent of these algorithms has problematized the very utility of theoretical analysis (Sculley et al., 2018), and the utility of statistical learning theory in particular. Namely, statistical learning theory seems simply not equipped to explain the generalization behaviour of machine learning algorithms like DNN’s (Belkin, 2021; Hardt and Recht, 2022; Berner et al., 2022), prompting a “quest for a new framework for a ‘theory of induction”’ (Belkin, 2021, p. 217). To put it bluntly: why a renewed sustained philosophical engagement with the framework of statistical learning theory, if this framework is starting to look like a thing of the past?
One answer is simply that it is still of philosophical interest if and how the standard framework already offers some justification for Occam’s razor. Curiously, statistical learning theory is largely left out the modern shift of the philosophical debate towards various frameworks in mathematical statistics (Sober, 2015, ch. 2; Baker, 2022, sect. 5), with Sober (2015, p. 140) perceiving statistical learning theory to be “dramatically” different from the “Bayesian and frequentist ideas” that have informed this debate so far. The current project thus fills a clear gap in the philosophical literature. Secondly, I may above have put things overly bluntly: it is not at all clear that core components of statistical learning theory will not continue to play an essential role in newer theory (cf. Bartlett et al., 2021). In any case, finally, the current project is a stepping stone towards the philosophical analysis of any new “framework for a theory of induction” in machine learning. Belkin indeed evokes “a very pure form of Occam’s razor” as the “guiding principle” in a new framework (2021, p. 218). To properly assess the role and justification of simplicity in such an emerging new framework, it will at the very least be helpful to actually have clarity on its role in the standard framework. The “generalization puzzle” (Berner et al., 2022, p. 25) that is now hotly debated in the machine learning community is indeed a modern reincarnation of exactly those traditional philosophical questions around the reliability of induction. Work like the current project can, I hope, provide a basis for philosophers to engage with this exciting but complex debate.
@@
The plan of this paper is as follows. In section 2, I present the framework of statistical learning theory and the main technical ingredients for the argument for Occam’s razor. These include the notions of empirical risk minimization, learnability, uniform convergence, and VC dimension, and the fundamental theorem that ties these notions together. In section 3, I argue that VC dimension is a robust notion of the simplicity of a hypothesis class. In section 4, I present the justificatory argument. This argument relies, as mentioned, on the model-relativity of learning-theoretic justification, and more generally on an epistemological perspective on machine learning methods that calls for this type of normative element. I conclude in section 5.
@@ 2.2. Hypothesis classes and learning methods
Empirical risk minimization for hypothesis class $\mathcal{H}$, write ${}_{\mathcal{H}}$, returns for each $S\in\mathcal{S}$ a hypothesis in $\operatorname*{arg\,min}_{h\in\mathcal{H}}L_{S}(h)$.
What makes a learning method like ERM for $\mathcal{H}$ a good method? Given the indicated goal of finding a relatively-low-risk hypothesis in $\mathcal{H}$, method ${}_{\mathcal{H}}$ can be called good if it has some sufficiently strong guarantee of attaining this goal.
@@ 2.3. Learnability
This guarantee is probably approximately correct (PAC) learnability [^6] This notion was formulated (with the additional component of computational complexity) by Valiant (1984), while the term “pac-learning” appears to be due to Angluin and Laird (1988). —or simply, learnability. Hypothesis class $\mathcal{H}$ is learnable by a learning method $A_{\mathcal{H}}$ if for any given accuracy $\epsilon$ and confidence $1-\delta$, there is a large enough sample size $m_{0}$ such that for any $m\geq m_{0}$, we have the following, no matter the true distribution $\mathcal{D}$. With probability at least $1-\delta$ over the possible size-$m$ samples $S^{m}$ drawn i.i.d. from $\mathcal{D}$, method $A_{\mathcal{H}}$, on receiving such a sample, returns a hypothesis with estimation error below $\epsilon$. To rephrase,
@@
Note that this guarantee (in particular, the minimum sample size $m_{\mathcal{H}}(\epsilon,\delta)$ for given $\epsilon$ and $\delta$) only depends on the hypothesis class $\mathcal{H}$. In line with the agnostic approach in statistical learning theory, it is a distribution-free guarantee: the sample size does not depend on the true distribution $\mathcal{D}$.
@@ 2.4. Uniform convergence and empirical risk minimization
The law of large numbers already tells us that, for any fixed hypothesis $h$, as we draw larger and larger samples $S^{m}$ i.i.d. from true distribution $\mathcal{D}$, the empirical error of $h$ on $S^{m}$ will in probability converge to its true risk. However, in our learning problem, we are not interested in fixing a particular hypothesis and estimating its true risk. We are interested in the performance of a learning algorithm, which, dependent on the data, can select different hypotheses. For this we need something stronger, namely a “uniform law of large numbers,” which bounds the difference between empirical errors and true risks of all hypotheses uniformly—simultaneously.
@@
On such a sample, all empirical errors give good indications of the true errors: “what you see is what you get” (“wysiwyg”, terminology Belkin, 2021). Now a hypothesis class has the uniform convergence property if there is a “wysiwyg” guarantee of drawing such representative samples. Precisely,
@@
Now recall that the method ${}_{\mathcal{H}}$ for given sample $S$ of length $m$ selects an $h$ that minimizes $L_{S}(h)$. Since $\epsilon_{\mathcal{H}}^{\mathrm{uc}}(m,\delta)$ is a constant term for fixed $m$ and $\delta$, method ${}_{\mathcal{H}}$ can be seen to explicitly minimize this upper bound
@@
on the true risk. Thus, given $\mathcal{H}$ satisfies uniform convergence, ${}_{\mathcal{H}}$ selects a hypothesis with the sharpest uniform upper bound on its true risk.
This minimization property, under the assumption of uniform convergence, allows us to derive that ${}_{\mathcal{H}}$ learns $\mathcal{H}$. Informally, [^7] See Shalev-Shwartz and Ben-David (2014, sect. 4.1, specifically lemma 4.2) for the (straightforward) formal derivation. if we have a guarantee that large enough samples are probably representative (uniform convergence), then in particular the lowest-empirical-error hypotheses (selected by ${}_{\mathcal{H}}$) probably have approximately lowest true risk, and so small estimation error (learnability).
@@ 2.6. Bringing it all together
Theorem 5 (Fundamental theorem of statistical learning theory [^10] In their pioneering work, Vapnik and Chervonenkis (1971) established the link between uniform convergence, “consistency” of ERM, and their notion of VC dimension, proving a generalization of the “fundamental theorem of mathematical statistics,” the Glivenko-Cantelli theorem of the uniform convergence of the empirical distribution function (see Devroye et al., 1996, ch. 12). The connection between VC theory and computational learning theory (in particular, Valiant’s notion of PAC learnability) was first spelled out by Blumer et al. (1986, 1989)., [^11] The most significant direction is from finite VC dimension to uniform convergence. Very roughly, the proof approach as presented by Shalev-Shwartz and Ben-David (2014) is as follows. The main component is a major generalization of the proof for finite classes (that essentially rests on two basic lemma’s, the union bound and Hoeffding’s inequality, see ibid., sect. 4.2) to a bound on the difference between true and empirical risk in terms of a growth function. This growth function tracks the “effective size” of the hypothesis class; for each $n$, it is the largest set $\mathcal{H}_{|X}$ for any $X$ of size $n$ (see footnote 8; ibid, def. 6.9). Next, we can show that for a class of VC dimension $d$, the growth function beyond $d$ grows, not exponentially, but only polynomially (Sauer’s lemma, see ibid., lmm. 6.10). Plugging this in into the bound, the result follows (see ibid., p. 51). Also see von Luxburg and Schölkopf (2011, sects. 5.1–5.6); Devroye et al. (1996, sect. 12.4).).
@@
  • $\mathcal{H}$ is learnable by ${}_{\mathcal{H}}$;
@@
In particular, if, and only if, hypothesis class $\mathcal{H}$ has finite VC dimension, we have a “wysiwyg” guarantee of a good indication of true risk (uniform convergence), and a method, ${}_{\mathcal{H}}$, that is a good method, in the sense of satisfying a guarantee of minimizing estimation error (learnability). We therefore have, for VC class $\mathcal{H}$, a certain justification for the ${}_{\mathcal{H}}$ method. [^12] In my presentation, I also follow Shalev-Shwartz and Ben-David (2014) in focusing on the epistemic end of learnability (minimizing estimation error). But another important epistemic end that is supported by the uniform convergence “wysiwyg” guarantees is model assessment, where we use the training error to assess whether the model (the learned hypothesis or indeed the hypothesis class) is actually good. (For instance, in the discussion of VC theory by Hastie et al., 2009, sect. 7.9, the emphasis is rather on this end.) I will continue to build the justificatory story on the inductive goal of learnability, but briefly return to this further end in footnote 39.
@@
At bottom, the fundamental theorem expresses a certain relation between four quantities (VC dimension, sample size, accuracy, and confidence), where, in particular, a lower VC dimension makes room for lower values of the other three quantities (meaning, for stronger bounds). Thus, a lower VC dimension of $\mathcal{H}$ goes with a better guarantee and therefore a stronger justification for the ${}_{\mathcal{H}}$ method.
@@ 3.1. Individual hypotheses and hypothesis classes
An influential example is the argument of Blumer et al. (1987) that “under very general assumptions, Occam’s Razor produces hypotheses that with high probability will be predictive of future observations” (ibid., p. 378). [^15] The relevant result is derived within Valiant’s PAC learning framework, and the requirement of computational efficiency in the definition and generalization guarantee of the relevant “Occam-algorithm” make it a bit more involved than the reasoning I discuss in this paper. For further expositions, see Anthony and Biggs (1992, pp. 59ff), Kearns and Vazirani (1994, ch. 2); and for a rebuttal of the argument, see Herrmann (2020). An earlier example still is Pearl (1978), who discusses the connection between simplicity and “credibility” of hypotheses via different notions of capacity and generalization success—including already VC dimension and uniform convergence (ibid., pp. 261ff). Pearl sets the stage as follows. We take some language $L$ with an interpretation function $I$ that maps sentences in the language to hypotheses (ibid., pp. 256f). Then we define some complexity measure on each sentence $t$, “which may represent either the syntactic aspect of the sentence $t$, or the work required for the computation of $I(t)$” (ibid., p. 257). [^16] The standard example is the two-symbol language of bits, where the complexity of a sentence (a bit string) is defined as its length. Further, “the complexity of a [hypothesis $h$] with respect to a language $L$ is defined as the complexity of the simplest sentence which represents that [hypothesis]” (ibid.). This allows us to take subsets of sufficiently simple hypotheses: a “complexity bounded sublanguage of $L$ is a sublanguage $L_{c}=(T_{c},I_{c})$ such that $T_{c}\subseteq T$, $I_{c}\subseteq I$ and $C(h)\leq c$ for all $h\in I_{c}$” (ibid., p. 258, slight change in notation).
@@
In particular, whereas classes of hypotheses of low representational complexity must be small, the converse does not hold. One can select small classes of (representationally) complex hypotheses, and the same capacity-based reasoning for good generalization still applies (cf. Mitchell, 1997, p. 65; Domingos, 1999, p. 410). [^18] This is the main critique of Herrmann (2020) of the argument of Blumer et al. Herrmann derives a parallel result for an “Anti-Occam algorithm” that selects small-cardinality classes of representationally complex hypotheses. A deeper problem still is that this notion of representational complexity depends on the presupposed formal language and definition of sentence complexity. For any hypothesis that is simple relative to one language, we can design a different language that renders it complex. [^19] This language-relativity in describing individual hypotheses also clearly arises in some presentations of the minimum description length (MDL) approach (e.g., Mitchell, 1997, sect. 6.6; Shalev-Shwartz and Ben-David, 2014, pp. 65f). However, these presentations paint a rather, well, simplistic picture of the approach: in “refined MDL,” the focus is on the design of “universal codes,” yielding again a robust notion of complexity of hypothesis classes (Grünwald, 2007) that plays a role very similar to capacity notions in statistical learning theory (ibid., sect. 17.10). In that sense representation length does not give us a robust or objective notion of simplicity of individual hypotheses. [^20] It is sometimes held that “idealized MDL” or Kolmogorov complexity can offer an objective notion of the representational complexity of individual objects (Li and Vitányi, 2008). See Sterkenburg (2016) for a critique of a suggested justification for Occam’s razor via this approach, and Sterkenburg (2018, sect. 5.2) for a critique of its promise of an objective notion of complexity. Given the bleak prospects for some general mathematical definition to settle what counts as simple for individual hypotheses, one might at this point be inclined to change tack and suggest that in practice, there is often no real problem. For many specific learning problems, we do appear to have clear intuitions about natural representations or parametrizations of hypotheses. In the standard curve-fitting problem (see Sober, 2015, pp. 88ff), for instance, where we seek to estimate a polynomial function, there exists a natural parametrization by degree. The linear functions of degree 1 are simpler than the quadratic functions of degree 2. Moreover, the class of all linear hypotheses is smaller (has lower capacity) than the strict superclass of quadratic hypotheses. Some conception of simplicity of hypotheses is here already taken for granted, which points to a natural ranking of hypothesis classes, and this ranking neatly aligns with their capacity. [^21] One might seek to base this conception on some formal definition of simplicity in terms of number of adjustable parameters, a line going back at least to Jeffreys (1939). But this still does not suddenly give us a robust and objective notion of simplicity of individual hypotheses: the “grue-like” problems of representation invariance remain (Priest, 1976). For a recent discussion and critique of defining simplicity by number of parameters, see Bonk (2023). If this is a common situation in practice, then, together with the formal connection between low capacity and generalization performance, we may have the basis for an explanation of why preferring simple hypotheses generally seems to be a good idea. But even if we accept as given, for many specific learning problems, a standard representation or parametrization of hypotheses, the formal connection between low capacity and generalization performance still falls short of constituting a justification for preferring simple hypotheses (for these specific learning problems). The issue remains that the theory does not enforce a connection between simple individual hypotheses (however specified) and classes of low capacity.

### v2

# Statistical learning theory and Occam’s razor: The core argument

@@ Abstract.
Statistical learning theory is often associated with the principle of Occam’s razor, which recommends a simplicity preference in inductive inference. This paper distills the core argument for simplicity obtainable from statistical learning theory, built on the theory’s central learning guarantee for the method of empirical risk minimization. This core “means-ends” argument is that a simpler hypothesis class or inductive model is better because it has better learning guarantees; however, these guarantees are model-relative and so the theoretical push towards simplicity is checked by our prior knowledge.
@@ 1. Introduction
Statistical learning theory is the standard framework for the mathematical analysis of machine learning methods (Vapnik, 2000; Shalev-Shwartz and Ben-David, 2014). The framework offers theoretical learning guarantees for certain learning methods, thus providing a basis for viewing such methods as good methods.
An old trope in machine learning, usually evoked under the label of Occam’s razor, is that a shared trait of good methods is a bias towards simplicity (Mitchell, 1997; Duda et al., 2001; Shalev-Shwartz and Ben-David, 2014; Goodfellow et al., 2016; Mohri et al., 2018; Alpaydin, 2020). Occam’s razor, understood as the principle that a simplicity preference is integral to good scientific or inductive reasoning, is also a long-standing topic of debate in the philosophy of science (Sober, 2015; Baker, 2022). The central question here is whether we actually have some epistemic justification for Occam’s razor. That is, we seek a rational reason for holding that a simplicity preference helps us attain desirable epistemic ends, like minimizing error.
@@
In this paper, I integrate the intuitions and arguments from both sides into a qualified epistemic justification for Occam’s razor from statistical learning theory. Importantly, the notion of simplicity in this justification pertains, not to individual classifiers or hypotheses, but to inductive models or classes of hypotheses. A further important component of my account is the relativity to such inductive models of the justification obtainable from theoretical learning guarantees, as highlighted by Sterkenburg and Grünwald (2021). This model-relativity of learning-theoretic justification aligns well, I argue, with a broad tradition in the philosophy of science which accepts the impossibility of absolute justification, and shifts attention to the project of how to rationally proceed from our current beliefs and assumptions. A final important characteristic of my account is the means-ends nature of the justificatory reasoning. In one line, the means-ends justificatory argument says that in order to have better model-relative learning guarantees, we need to codify our assumptions in the form of a simpler inductive model.
@@
This argument is based on the first of the two “inductive principles” (Vapnik, 2000) central to statistical learning theory, namely the method of empirical risk minimization. I think this is the “core argument” for simplicity in statistical learning theory, which further underpins the method of structural risk minimization and its characteristic simplicity preference. I sketch this at the end of the paper.
@@
The plan of the paper is as follows. In section 2, I present the framework of statistical learning theory and the main technical ingredients for the core argument for Occam’s razor. These include the notions of empirical risk minimization, learnability, uniform convergence, and VC dimension, and the fundamental theorem that ties these notions together. In section 3, I argue that VC dimension is a robust notion of the simplicity of a hypothesis class. In section 4, I discuss the theoretical justification for empirical risk minimization and particularly its model-relative nature; and I show how all of the previous comes together into a justificatory argument for simplicity. I conclude in section 5.
Motivation
Before starting, there is a worry about the paper’s general project that I should acknowledge. This worry is that the project engages with a debate of a bygone era. It has been well over a decade since Harman and Kulkarni (2007) initiated a small wave of philosophical interest in statistical learning theory, and Steel (2011, p. 860) concluded that the theory is “worthy of further sustained interest from philosophers of science.” This sustained interest has not exactly materialized, while the landscape of machine learning has altered significantly. Especially the advent of deep neural networks (DNN’s) has caused a shift in what seem the more pertinent epistemological issues: from the traditional questions around the reliability of inductive inference to questions around interpretability and explainability (Beisbart and Räz, 2022). Moreover, the advent of these algorithms has problematized the very utility of statistical learning theory. The theory seems simply not equipped to explain the generalization behaviour of learning methods like DNN’s (Belkin, 2021; Hardt and Recht, 2022; Berner et al., 2022), prompting a “quest for a new framework for a ‘theory of induction”’ (Belkin, 2021, p. 217). Putting it bluntly: why a renewed philosophical engagement with the framework of statistical learning theory, if this framework is starting to look like a thing of the past?
One plain answer is that it is still of interest if and how the standard framework already offers justification for Occam’s razor. Curiously, statistical learning theory is largely left out the modern shift of the philosophical debate towards various frameworks in mathematical statistics (Sober, 2015, ch. 2; Baker, 2022, sect. 5), with Sober (2015, p. 140, fn. 61) perceiving statistical learning theory to be “dramatically” different from the “Bayesian and frequentist ideas” that have informed this debate so far. The current project thus fills a gap in the philosophical literature. Secondly, I may above have put things overly bluntly: it is not at all clear that core components of statistical learning theory will not continue to play an essential role in newer theory (cf. Bartlett et al., 2021). In any case, finally, the current project is a stepping stone towards the philosophical analysis of any new “framework for a theory of induction” in machine learning. Belkin indeed evokes “a very pure form of Occam’s razor” as the “guiding principle” in a new framework (2021, p. 218). To assess the role and justification of simplicity in such an emerging new framework, it will at the very least be helpful to actually have clarity on its role in the standard framework. The “generalization puzzle” (Berner et al., 2022, p. 25) that is now hotly debated in the machine learning community is indeed a modern reincarnation of exactly those traditional philosophical questions around the reliability of induction. Work like the current project can, I hope, offer a starting point for philosophers to engage with this exciting but complex debate.
@@ 2.2. Hypothesis classes and learning methods
Empirical risk minimization for hypothesis class $\mathcal{H}$, write ERMH, returns for each $S\in\mathcal{S}$ a hypothesis in $\operatorname*{arg\,min}_{h\in\mathcal{H}}L_{S}(h)$.
What makes a learning method like ERM for $\mathcal{H}$ a good method? Given the indicated goal of finding a relatively-low-risk hypothesis in $\mathcal{H}$, method ERMH can be called good if it has some sufficiently strong guarantee of attaining this goal.
@@ 2.3. Learnability
This guarantee is probably approximately correct (PAC) learnability [^6] This notion was formulated (with the additional component of computational complexity) by Valiant (1984), while the term “pac-learning” appears to be due to Angluin and Laird (1988). —or simply, learnability. Hypothesis class $\mathcal{H}$ is learnable by a learning method $A_{\mathcal{H}}$ if for any given inaccuracy $\epsilon$ and confidence $1-\delta$, there is a large enough sample size $m_{0}$ such that for any $m\geq m_{0}$, we have the following, no matter the true distribution $\mathcal{D}$. With probability at least $1-\delta$ over the possible size-$m$ samples $S^{m}$ drawn i.i.d. from $\mathcal{D}$, method $A_{\mathcal{H}}$, on receiving such a sample, returns a hypothesis with estimation error below $\epsilon$. To rephrase,
@@
Note that this guarantee (in particular, the minimum sample size $m_{\mathcal{H}}(\epsilon,\delta)$ for given $\epsilon$ and $\delta$) only depends on the hypothesis class $\mathcal{H}$. In line with the agnostic approach of statistical learning theory, it is a distribution-free guarantee: the sample size does not depend on the true distribution $\mathcal{D}$.
@@ 2.4. Uniform convergence and empirical risk minimization
The law of large numbers already tells us that, for any fixed hypothesis $h$, as we draw larger and larger samples $S^{m}$ i.i.d. from true distribution $\mathcal{D}$, the empirical error of $h$ on $S^{m}$ will in probability converge to its true risk. However, in our learning problem, we are not interested in fixing a particular hypothesis and estimating its true risk. We are interested in the performance of a learning algorithm, which, depending on the data, can select different hypotheses. For this we need something stronger, namely a “uniform law of large numbers,” which bounds the difference between empirical errors and true risks of all hypotheses uniformly—simultaneously.
@@
On such a sample, all empirical errors give good indications of the true errors: “what you see is what you get” (“wysiwyg”, terminology from Belkin, 2021). Now a hypothesis class has the uniform convergence property if there is a “wysiwyg” guarantee of drawing such representative samples. Precisely,
@@
Now recall that the method ERMH for given sample $S$ of length $m$ selects an $h$ that minimizes $L_{S}(h)$. Since $\epsilon_{\mathcal{H}}^{\mathrm{uc}}(m,\delta)$ is a constant term for fixed $m$ and $\delta$, method ERMH can be seen to explicitly minimize this upper bound
@@
on the true risk. Thus, given $\mathcal{H}$ satisfies uniform convergence, ERMH selects a hypothesis with the sharpest uniform upper bound on its true risk.
This minimization property, under the assumption of uniform convergence, allows us to derive that ERMH learns $\mathcal{H}$. Informally, [^7] See Shalev-Shwartz and Ben-David (2014, sect. 4.1, specifically lemma 4.2) for the (straightforward) formal derivation. if we have a guarantee that large enough samples are probably representative (uniform convergence), then in particular the lowest-empirical-error hypotheses (selected by ERMH) probably have approximately lowest true risk, and so small estimation error (learnability).
@@ 2.6. Bringing it all together
Theorem 5 (Fundamental theorem of statistical learning theory [^10] In their pioneering work, Vapnik and Chervonenkis (1971) established the link between uniform convergence, “consistency” of ERM, and their notion of VC dimension, proving a generalization of the “fundamental theorem of mathematical statistics,” the Glivenko-Cantelli theorem of the uniform convergence of the empirical distribution function (see Devroye et al., 1996, ch. 12). The connection between VC theory and computational learning theory (in particular, Valiant’s notion of PAC learnability) was first spelled out by Blumer et al. (1986, 1989).).
@@
  • $\mathcal{H}$ is learnable by ERMH;
@@
In particular, if, and only if, hypothesis class $\mathcal{H}$ has finite VC dimension, we have a “wysiwyg” guarantee of a good indication of true risk (uniform convergence), and a method, ERMH, that is a good method, in the sense of satisfying a guarantee of minimizing estimation error (learnability). We therefore have, for VC class $\mathcal{H}$, a certain justification for the ERMH method. [^11] In my presentation, I also follow Shalev-Shwartz and Ben-David (2014) in focusing on the epistemic end of learnability (minimizing estimation error). I only note here that another important epistemic end that is supported by the uniform convergence “wysiwyg” guarantees is model assessment, where we use the training error to assess whether the model (the learned hypothesis or indeed the hypothesis class) is actually good. (For instance, in the discussion of VC theory by Hastie et al., 2009, sect. 7.9, the emphasis is rather on this end.)
@@
At bottom, the fundamental theorem expresses a certain relation between four quantities (VC dimension, sample size, accuracy, and confidence), where, in particular, a lower VC dimension makes room for lower values of the other three quantities (meaning, for stronger bounds). Thus, a lower VC dimension of $\mathcal{H}$ goes with a better guarantee and therefore a stronger justification for the ERMH method.
@@ 3.1. Individual hypotheses and hypothesis classes
An influential example is the argument of Blumer et al. (1987) that “under very general assumptions, Occam’s Razor produces hypotheses that with high probability will be predictive of future observations” (ibid., p. 378). [^14] The relevant result is derived within Valiant’s PAC learning framework, and the requirement of computational efficiency in the definition and generalization guarantee of the relevant “Occam-algorithm” make it a bit more involved than the reasoning I discuss in this paper. For further expositions, see Anthony and Biggs (1992, pp. 59ff), Kearns and Vazirani (1994, ch. 2); and for a rebuttal of the argument, see Herrmann (2020). An earlier example still is Pearl (1978), who discusses the connection between simplicity and “credibility” of hypotheses via different notions of capacity and generalization success—including already VC dimension and uniform convergence (ibid., pp. 261ff).
Pearl sets the stage as follows. We take some language $L$ with an interpretation function $I$ that maps sentences in the language to hypotheses (ibid., pp. 256f). Then we define some complexity measure on each sentence $t$, “which may represent either the syntactic aspect of the sentence $t$, or the work required for the computation of $I(t)$” (ibid., p. 257). [^15] The standard example is the two-symbol language of bits, where the complexity of a sentence (a bit string) is defined as its length. Further, “the complexity of a [hypothesis $h$] with respect to a language $L$ is defined as the complexity of the simplest sentence which represents that [hypothesis]” (ibid.). This allows us to take subsets of sufficiently simple hypotheses: a “complexity bounded sublanguage of $L$ is a sublanguage $L_{c}=(T_{c},I_{c})$ such that $T_{c}\subseteq T$, $I_{c}\subseteq I$ and $C(h)\leq c$ for all $h\in I_{c}$” (ibid., p. 258, slight change in notation).
@@
In particular, whereas classes of hypotheses of low representational complexity must be small, the converse does not hold. One can select small classes of (representationally) complex hypotheses, and the same capacity-based reasoning for good generalization still applies (cf. Mitchell, 1997, p. 65; Domingos, 1999, p. 410). [^17] This is the main critique of Herrmann (2020) of the argument of Blumer et al. Herrmann derives a parallel result for an “Anti-Occam algorithm” that selects small-cardinality classes of representationally complex hypotheses.
A deeper problem still is that this notion of representational complexity depends on the presupposed formal language and definition of sentence complexity. For any hypothesis that is simple relative to one language, we can design a different language that renders it complex. [^18] This language-relativity in describing individual hypotheses also clearly arises in some presentations of the minimum description length (MDL) approach (e.g., Mitchell, 1997, sect. 6.6; Shalev-Shwartz and Ben-David, 2014, pp. 65f). However, these presentations paint a rather, well, simplistic picture of the approach: in “refined MDL,” the focus is on the design of “universal codes,” yielding again a robust notion of complexity of hypothesis classes (Grünwald, 2007) that plays a role very similar to capacity notions in statistical learning theory (ibid., sect. 17.10). In that sense representation length does not give us a robust or objective notion of simplicity of individual hypotheses. [^19] It is sometimes held that “idealized MDL” or Kolmogorov complexity can offer an objective notion of the representational complexity of individual objects (Li and Vitányi, 2008). See Sterkenburg (2016) for a critique of a suggested justification for Occam’s razor via this approach, and Sterkenburg (2018, sect. 5.2) for a critique of its promise of an objective notion of complexity.
@@
Given the bleak prospects for some general mathematical definition to settle what counts as simple for individual hypotheses, one might at this point be inclined to change tack and suggest that in practice, there is often no real problem. For many specific learning problems, we do appear to have clear intuitions about natural representations or parametrizations of hypotheses. In the standard curve-fitting problem (see, e.g., Sober, 2015, pp. 88ff), where we seek to estimate a polynomial function, there exists a natural parametrization by degree. [^20] Curve-fitting can be cast as a problem in binary classification by treating the curves as hypotheses separating instances with the one label from instances with the other. The linear functions of degree 1 are simpler than the quadratic functions of degree 2. Moreover, the class of all linear hypotheses is smaller (has lower capacity) than the strict superclass of quadratic hypotheses. Some conception of simplicity of hypotheses is here already taken for granted, which points to a natural ranking of hypothesis classes, and this ranking neatly aligns with their capacity. [^21] One might seek to base this conception on some formal definition of simplicity in terms of number of adjustable parameters, a line going back at least to Jeffreys (1939). But this still does not suddenly give us a robust and objective notion of simplicity of individual hypotheses: the “grue-like” problems of representation invariance remain (Priest, 1976). For a recent discussion and critique of defining simplicity by number of parameters, see Bonk (2023).
If this is a common situation in practice, then, together with the formal connection between low capacity and generalization performance, we may have the basis for an explanation of why preferring simple hypotheses generally seems to be a good idea. But even if we accept as given, for many specific learning problems, a standard representation or parametrization of hypotheses, the formal connection between low capacity and generalization performance still falls short of constituting a justification for preferring simple hypotheses (for these specific learning problems). The issue remains that the theory does not enforce a connection between simple individual hypotheses (however specified) and classes of low capacity.